# Efficient and Context-Aware Label Propagation for Zero-/Few-Shot Training-Free Adaptation of Vision-Language Model

### v1

@@ Abstract
Vision-language models (VLMs) have revolutionized machine learning by leveraging large pre-trained models to tackle various downstream tasks. Despite improvements in label, training, and data efficiency, many state-of-the-art VLMs still require task-specific hyperparameter tuning and fail to fully exploit test samples. To overcome these challenges, we propose a graph-based approach for label-efficient adaptation and inference. Our method dynamically constructs a graph over text prompts, few-shot examples, and test samples, using label propagation for inference without task-specific tuning. Unlike existing zero-shot label propagation techniques, our approach requires no additional unlabeled support set and effectively leverages the test sample manifold through dynamic graph expansion. We further introduce a context-aware feature re-weighting mechanism to improve task adaptation accuracy. Additionally, our method supports efficient graph expansion, enabling real-time inductive inference. Extensive evaluations on downstream tasks, such as fine-grained categorization and out-of-distribution generalization, demonstrate the effectiveness of our approach.
@@ 1 Introduction
Emerging foundation models have transformed the traditional paradigm of machine learning model development (Jia et al., 2021; Kim et al., 2021; Li et al., 2022). By pre-training vision-language models (VLMs) on a massive scale of image-language pairs, these models have demonstrated strong inference capabilities across a wide range of downstream tasks, all while reducing the need for extensive data collection and labeling (Zhou et al., 2022b; Zhang et al., 2022; Shu et al., 2022).
Recent research on adapting pre-trained VLMs to downstream tasks has primarily focused on improving label efficiency, training efficiency, and data efficiency. DMN (Zhang et al., 2024b) represents a comprehensive approach that addresses all three dimensions. Specifically, DMN leverages text prompts, test samples, and few-shot samples to construct a three-branch classifier, with final predictions being the fusion of the individual branches. While DMN achieves state-of-the-art performance on fine-grained categorization tasks, it requires tuning a task-specific coefficient to fuse predictions, often using the test set for hyper-parameter tuning in the absence of a dedicated validation set. Additionally, DMN introduces a memory bank of test samples to synthesize an adaptive classifier for each sample. However, we hypothesize that these test samples can be used more effectively—not just for classifier synthesis but also to better capture the data manifold for transductive inference (Joachims, 2003), particularly when labeled samples are limited.
To reduce the need for task-specific hyper-parameter tuning and better utilize test samples, we propose a graph-based adaptation and inference method for downstream tasks. Our approach dynamically constructs a graph based on available few-shot samples (if applicable), test samples, and text prompts linked to semantic labels. This graph captures the intrinsic data manifold, and we use label propagation (Zhu & Ghahramani, 2002) for inference. Unlike DMN, this method eliminates the need for task-specific hyper-parameter tuning and enhances the use of information embedded in the unlabeled test samples.
@@
Figure 1: Illustration of the overall framework of ECALP. We use the text prompts, few-shot samples, and testing samples to build the graph. The graph is dynamically expanded upon seeing a new sample with context-aware re-weighted edge weights. An iterative solution is adopted to predict labels for all testing samples.
In this work, we introduce a holistic label propagation approach to label-efficient adaptation of VLMs with an overview presented in Fig. 1. To address the challenge of computational efficiency, we employ an iterative solution to label propagation instead of the closed-form solution, which could benefit from incremental label propagation and label reset. Furthermore, we employ a context-aware feature channel re-weighting to better adapt to downstream tasks. We require only the text prompts and/or few-shot samples to provide contextual information for feature re-weighting. We also present an efficient graph expansion mechanism to allow inductive inference on a stream of test samples, without requiring all test data for transductive labeling. Our proposed method is referred to as Efficient and Context-Aware Label Propagation (ECALP). Experiments are conducted on a wide range of downstream tasks, including fine-grained categorization, distribution shift and few-shot categorization. We summarize the contributions of this work as follows:
  • We propose a unified label-efficient adaptation of vision-language models from a label propagation perspective. Compared with state-of-the-art methods, the proposed method achieves higher inference speed and fixed hyper-parameters.
  • We account for the diversity of information captured by the vision encoder and propose to re-weight the vision feature channels using the statistics of the text embeddings and/or few-shot samples’ features from the downstream task.
@@ 2.1 Zero/Few-Shot Adaptation of Vision-Language Model
Vision-language models (VLMs)(Zhang et al., 2024a; Jia et al., 2021; Li et al., 2022; 2023) have achieved remarkable success in open-vocabulary classification tasks. However, their performance significantly declines in fine-grained scenarios(Zhao et al., 2017) or under distribution shifts (Hendrycks & Dietterich, 2019), highlighting the limitations of large-scale pre-training. To address these challenges, two primary strategies have emerged to enhance VLM adaptability in zero-shot and few-shot settings: prompt tuning and adapter tuning.
Prompt tuning (Yoon et al., 2024; Feng et al., 2023), inspired by techniques in language models (Brown, 2020), improves performance by optimizing learnable input prompts. For instance, CoOp (Zhou et al., 2022b) and CoCoOp (Zhou et al., 2022a) replace static text prompts with learnable word vectors, fine-tuned via few-shot classification loss. TPT (Shu et al., 2022) and SwapPrompt (Ma et al., 2024) extend this concept to zero-shot settings, employing contrastive learning with augmented image inputs to generate pseudo-labels. Although effective, prompt tuning requires full backpropagation through the computational graph, which results in high computational costs.
Adapter tuning (Zhu et al., 2023; Wang et al., 2024), as demonstrated by Clip-Adapter (Gao et al., 2024), adapts image or text features using lightweight adapters, thereby avoiding full backpropagation through complex encoders. Recently, training-free adapter approaches, such as Tip-Adapter (Zhang et al., 2022), have gained traction by leveraging few-shot image features as prototypes, removing the need for learned adapter weights. DMN (Zhang et al., 2024b) and TDA (Karmanov et al., 2024) extend these methods to zero-shot scenarios by building memory banks from test data and pseudo-labels. However, these approaches often introduce additional costs due to image augmentation and repeated sample processing. Moreover, their reliance on complex architectures and extensive hyperparameter tuning limits their practicality in real-world applications.
In this work, we propose a unified, training-free adaptation framework for VLMs based on label propagation. Our method exploits the manifold structure of test data, eliminating the need for augmentation and exhaustive hyper-parameter searches, offering a more efficient and practical solution compared to previous approaches.
@@ 2.2 Label Propagation
Label propagation (LP) is a well-established method widely used for label-efficient learning (Zhu et al., 2003; Iscen et al., 2019; Xu & Lee, 2020; Zhu & Koniusz, 2023). LP operates under the assumption that labels vary smoothly across a graph, with adjacent nodes likely sharing the same label. In transductive learning, where all test samples are visible during inference, LP propagates labels from labeled nodes to unlabeled ones. It has been particularly effective in retrieval tasks, where a new query is appended to the graph, and its affinity is propagated to prototype nodes (Iscen et al., 2017). Recent works have demonstrated the benefits of LP in adapting VLMs for downstream tasks (Hu et al., 2024; Kalantidis et al., 2024), where a graph is constructed from downstream test samples, and labels are propagated from text prompt prototypes to test samples using a closed-form LP solution. Despite its strengths, this approach faces several challenges, as previously discussed. In response, we propose a holistic solution for label-efficient adaptation of VLMs, built upon a label propagation framework.
@@ 3.2 Revisiting Transductive Label Propagation
$Y^{t+1}=\alpha\tilde{W}Y^{t}+(1-\alpha)Y^{0}\quad\Rightarrow\quad Y^{\infty}=(\mathbf{1}-\alpha\tilde{W})^{-1}Y^{0}$ (1)
@@
Efficient Label Propagation via Iterative Solution: The closed-form solution to label propagation was adopted by Kalantidis et al. (2024). However, we argue that this closed-form solution is sub-optimal in practice for two reasons. First, the closed-form solution requires solving a linear system using the conjugate gradient method (Grady, 2006). This is an expensive step that prohibits efficient inference. Furthermore, the closed-form solution is the converged solution to the iterative method. Each iteration step involves propagating labels between all connected nodes, including the inferred label information from testing samples back to the text prototypes and few-shot labeled samples, which is undesirable. For this reason, we employ the iterative solution and we reset the labels for textual prototypes and labeled samples after each iteration, i.e. $Y^{t+1}_{p}=Y^{0}_{p},Y_{l}^{t+1}=Y_{l}^{0}$.
@@ 3.3 Dynamic Graph Expansion for Inductive Inference
Static Graph Construction: We first introduce the way to construct a static graph when the whole set of unlabeled testing samples $\mathcal{D}_{u}$, textual prototypes $\mathcal{P}$, and few-shot samples (optional) $\mathcal{D}_{l}$ are available. Nodes are denoted as $\mathcal{V}=\{u_{1},...,u_{N_{u}},p_{1},...,p_{N_{p}},\l_{1},...,l_{N_{l}}\}$. We write the adjacency matrix $W$ blockwise in Eq. 3 where $W_{u}\in\mathbbm{R}^{N_{u}\times N_{u}}$, $W_{p}\in\mathbbm{R}^{N_{p}\times N_{p}}$, $W_{l}\in\mathbbm{R}^{N_{l}\times N_{l}}$, $W_{up}\in\mathbbm{R}^{N_{u}\times N_{p}}$, $W_{ul}\in\mathbbm{R}^{N_{u}\times N_{l}}$, $W_{pl}\in\mathbbm{R}^{N_{p}\times N_{l}}$. We do not connect any nodes between textual prototypes and few-shot samples because we are only interested in inferring the labels of unlabeled testing samples. This results in $W_{p}=\mathbf{0}$, $W_{l}=\mathbf{0}$, $W_{pl}=\mathbf{0}$.
@@
For label propagation, we sparsify $W$ sparse for improved computation cost and robustness. Observations made in (Zhu et al., 2023; Kalantidis et al., 2024) suggest that cosine similarities between inputs from different modalities can vary significantly. For example, the cosine similarity between the features of two images is generally significantly lower than between an image feature and a text embedding. Therefore, we search for the K-nearest neighbors within each modality separately as follows, with $\text{NN}k_{\mathcal{D}}$ referring to the $k$ nearest neighbor search within nodes from set $\mathcal{D}$.
@@
Dynamic Graph Expansion: The static graph construction assumes unlabeled testing samples are all available at once. This assumption prohibits inductive inference where testing samples arrive in a stream fashion. Previous work (Kalantidis et al., 2024) proposed an inductive inference approach by exploiting an external training dataset to build the manifold. With access to such an external dataset, a naive way to support inductive inference is to build a new graph from scratch upon seeing new testing data, which results in high computation complexity. For example, with a total number of $N_{v}=N_{p}+N_{l}+N_{u}$ nodes, each with $d$ channels, the computational complexity is $O(d\cdot N_{v}^{3}+N_{v}^{2}\log N_{v})$, which prohibits realistic inference on large scale downstream task. To enable inductive inference while exploiting the testing data manifold, we reduce the complexity to $O(d\cdot N_{v}^{2})$ by a dynamic graph expansion mechanism. Specifically, with an existing adjacency matrix $W$, we use the new testing sample $u_{N_{u}+1}$ to query all existing nodes and replace the weakest connections with connections to the new testing sample following the update rule in Eq. 5. The update rule will insert a new testing sample into the existing graph without re-calculating the K nearest neighbor of the existing nodes.
@@
Incremental Label Propagation: With the dynamically-expanded graph, we can perform label propagation incrementally. Specifically, we re-use the pseudo labels from the previous iteration to speed up the convergence of the label propagation process. After inference for each testing sample, the pseudo labels are attenuated for label propagation for the next testing sample. We follow the rules in Eq. 6 to produce the incremental labels $Y_{u}^{0}$ for the observed testing samples for inference on future testing samples, where $Y^{T}_{u}$ refers to the labels after $T$ steps of label propagation.
@@ 3.4 Context-aware Edge Re-Weighting
The graph construction process depends on a properly chosen distance metric. Measuring the distance between samples based on the features from the vision encoder is subject to the biases learnt by the vision encoder during large-scale pre-training. For example, it is known that VLMs encode features that capture all visual instances in the image, e.g. objects, background, style, etc. For this reason, directly measuring the cosine similarity between the raw features encoded by the VLM is subject to irrelevant information for the downstream task. For example, adapting the VLM to a downstream task of car model classification requires less attention to certain aspects like background or color information. Motivated by this, we propose to re-weight the importance of VLM vision-encoded features for calculating similarity. Specifically, we first calculate the statistics of textual prompts feature channels as follows, where $c$ is the feature channel index.
@@
A higher variance in $c$-th channel suggests the higher ‘discriminability’ of the channel, i.e. the more capable it is to discriminate between text prompts. Therefore, we increase the importance of these channels with higher variance. Additionally, we calculate the same statistics on the few-shot samples, resulting in $\mu_{c}^{l}$ and $\sigma_{c}^{l}$. In contrast to the context of textual prompts, we argue that high-variance feature channels in the context of few-shot samples highlight the intra-class variance and that this should be suppressed when calibrating the similarity. Therefore, we use the reciprocal of the variance for re-weighting graph edge similarity. The new edge weights are presented in Eq. 8:

### v2

@@ Abstract
Vision-language models (VLMs) have revolutionized machine learning by leveraging large pre-trained models to tackle various downstream tasks. Although label, training, and data efficiency have improved, many state-of-the-art VLMs still require task-specific hyperparameter tuning and fail to fully exploit test samples. To overcome these challenges, we propose a graph-based approach for label-efficient adaptation and inference. Our method dynamically constructs a graph over text prompts, few-shot examples, and test samples, using label propagation for inference without task-specific tuning. Unlike existing zero-shot label propagation techniques, our approach requires no additional unlabeled support set and effectively leverages the test sample manifold through dynamic graph expansion. We further introduce a context-aware feature re-weighting mechanism to improve task adaptation accuracy. Additionally, our method supports efficient graph expansion, enabling real-time inductive inference. Extensive evaluations on downstream tasks, such as fine-grained categorization and out-of-distribution generalization, demonstrate the effectiveness of our approach. The source code is available at https://github.com/Yushu-Li/ECALP.
**footnotetext: Equal contribution. Correspondence to Xun Xu: xu_xun@i2r.a-star.edu.sg. This work was done during Yushu Li and Yongyi Su’s visit to I2R.
@@ 1 Introduction
Emerging foundation models have transformed the traditional paradigm of machine learning model development (Jia et al., 2021; Kim et al., 2021; Li et al., 2022). By pre-training vision-language models (VLMs) on image-language pairs at a massive scale, these models have demonstrated strong inference capabilities across a wide range of downstream tasks, all while reducing the need for extensive data collection and labeling (Zhou et al., 2022b; Zhang et al., 2022; Shu et al., 2022).
Recent research on adapting pre-trained VLMs to downstream tasks has primarily focused on improving label efficiency, training efficiency, and data efficiency. DMN (Zhang et al., 2024b) represents a comprehensive approach that addresses all three dimensions. Specifically, DMN leverages text prompts, test samples, and few-shot samples to construct a three-branch classifier, with final predictions being the fusion of the individual branches. While DMN achieves state-of-the-art performance on fine-grained categorization tasks, it requires tuning a task-specific coefficient to fuse predictions, often using the test set for hyperparameter tuning in the absence of a dedicated validation set. Additionally, DMN introduces a memory bank of test samples to synthesize an adaptive classifier for each sample. However, we hypothesize that these test samples can be used more effectively—not just for classifier synthesis but also to better capture the data manifold for transductive inference (Joachims, 2003), particularly when labeled samples are limited.
To reduce the need for task-specific hyperparameter tuning and better utilize test samples, we propose a graph-based adaptation and inference method for downstream tasks. Our approach dynamically constructs a graph based on available few-shot samples (if applicable), test samples, and text prompts linked to semantic labels. This graph captures the intrinsic data manifold, and we use label propagation (Zhu & Ghahramani, 2002) for inference. Unlike DMN, this method eliminates the need for task-specific hyperparameter tuning and enhances the use of information embedded in the unlabeled test samples.
@@
Figure 1: Illustration of the overall framework of ECALP. We use the text prompts, few-shot samples, and test samples to build the graph. The graph is dynamically expanded upon seeing a new sample with context-aware re-weighted edge weights. An iterative solution is adopted to predict labels for all test samples.
In this work, we introduce a holistic label propagation approach to label-efficient adaptation of VLMs with an overview presented in Fig. 1. To address the challenge of computational efficiency, we employ an iterative solution to label propagation instead of the closed-form solution, which could benefit from incremental label propagation and label reset. Furthermore, we employ a context-aware feature dimension re-weighting to better adapt to downstream tasks. We require only the text prompts and/or few-shot samples to provide contextual information for feature re-weighting. We also present an efficient graph expansion mechanism to allow inductive inference on a stream of test samples, without requiring all test data for transductive labeling. Our proposed method is referred to as Efficient and Context-Aware Label Propagation (ECALP). Experiments are conducted on a wide range of downstream tasks, including fine-grained categorization, distribution shift and few-shot categorization. We summarize the contributions of this work as follows:
  • We propose a unified label-efficient adaptation of vision-language models from a label propagation perspective. Compared with state-of-the-art methods, the proposed method achieves higher inference speed and fixed hyperparameters.
  • We account for the diversity of information captured by the vision encoder and propose to re-weight the vision feature dimensions using the statistics of the text embeddings and/or few-shot samples’ features from the downstream task.
@@ 2.1 Zero/Few-Shot Adaptation of Vision-Language Model
Vision-language models (VLMs) (Zhang et al., 2024a; Jia et al., 2021; Li et al., 2022; 2023) have achieved remarkable success in open-vocabulary classification tasks. However, their performance significantly declines in fine-grained scenarios (Zhao et al., 2017) or under distribution shifts (Hendrycks & Dietterich, 2019), highlighting the limitations of large-scale pre-training. To address these challenges, two primary strategies have been developed to improve VLM adaptability in zero-shot and few-shot settings: prompt tuning and adapter tuning.
Prompt tuning (Yoon et al., 2024; Feng et al., 2023), inspired by techniques in language models (Brown et al., 2020), improves performance by optimizing learnable input prompts. For instance, CoOp (Zhou et al., 2022b) and CoCoOp (Zhou et al., 2022a) replace static text prompts with learnable word vectors, fine-tuned via few-shot classification loss. TPT (Shu et al., 2022) and SwapPrompt (Ma et al., 2024) extend this concept to zero-shot settings, employing contrastive learning on augmented image inputs to generate pseudo-labels. Although effective, prompt tuning requires full backpropagation through the computational graph, which results in high computational costs.
Adapter tuning (Zhu et al., 2023; Wang et al., 2024), exemplified by Clip-Adapter (Gao et al., 2024), adapts image or text features using lightweight adapters, thereby avoiding full backpropagation through complex encoders. Recently, training-free adapter approaches, such as Tip-Adapter (Zhang et al., 2022), have gained traction by leveraging few-shot image features as prototypes, removing the need for learned adapter weights. DMN (Zhang et al., 2024b) and TDA (Karmanov et al., 2024) extend these methods to zero-shot scenarios by building memory banks from test data and pseudo-labels. However, these approaches often introduce additional costs due to image augmentation and repeated sample processing. Moreover, their reliance on complex architectures and extensive hyperparameter tuning limits their practicality in real-world applications.
In this work, we propose a unified, training-free adaptation framework for VLMs based on label propagation. Our method exploits the manifold structure of test data, eliminating the need for augmentation and exhaustive hyperparameter searches, offering a more efficient and practical solution compared to previous approaches.
@@ 2.2 Label Propagation
Label propagation (LP) is a well-established method widely used for label-efficient learning (Zhu et al., 2003; Iscen et al., 2019; Xu & Lee, 2020; Zhu & Koniusz, 2023). LP operates under the assumption that labels vary smoothly across a graph, with adjacent nodes likely sharing the same label. In transductive learning, where all test samples are visible during inference, LP propagates labels from labeled nodes to unlabeled ones. It has been particularly effective in retrieval tasks, where a new query is appended to the graph, and its affinity is propagated to prototype nodes (Iscen et al., 2017). Recent studies have demonstrated the benefits of LP in adapting VLMs to downstream tasks (Hu et al., 2024; Kalantidis et al., 2024), where a graph is constructed from downstream test samples, and labels are propagated from text prompt prototypes to test samples using a closed-form LP solution. Despite its strengths, this approach faces several challenges, as previously discussed. In response, we propose a holistic solution for label-efficient adaptation of VLMs, built upon a label propagation framework.
@@ 3.2 Revisiting Transductive Label Propagation
$Y^{t+1}=\alpha\tilde{W}Y^{t}+(1-\alpha)Y^{0}\quad\Rightarrow\quad Y^{\infty}=(\mathbb{I}-\alpha\tilde{W})^{-1}Y^{0}$ (1)
@@
Efficient Label Propagation via Iterative Solution: The closed-form solution to label propagation was adopted by Kalantidis et al. (2024). However, we argue that this closed-form solution is sub-optimal in practice for two reasons. First, the closed-form solution requires solving a linear system using the conjugate gradient method (Grady, 2006). This is an expensive step that prohibits efficient inference. Furthermore, the closed-form solution is the converged solution to the iterative method. Each iteration step involves propagating labels between all connected nodes, including the inferred label information from test samples back to the text prototypes and few-shot labeled samples, which is undesirable. For this reason, we employ the iterative solution and we reset the labels for textual prototypes and labeled samples after each iteration, i.e. $Y^{t+1}_{p}=Y^{0}_{p},Y_{l}^{t+1}=Y_{l}^{0}$.
@@ 3.3 Dynamic Graph Expansion for Inductive Inference
Static Graph Construction: We first introduce the way to construct a static graph when the whole set of unlabeled test samples $\mathcal{D}_{u}$, textual prototypes $\mathcal{P}$, and few-shot samples (optional) $\mathcal{D}_{l}$ are available. Nodes are denoted as $\mathcal{V}=\{u_{1},...,u_{N_{u}},p_{1},...,p_{N_{p}},\l_{1},...,l_{N_{l}}\}$. We write the adjacency matrix $W$ blockwise in Eq. 3 where $W_{u}\in\mathbbm{R}^{N_{u}\times N_{u}}$, $W_{p}\in\mathbbm{R}^{N_{p}\times N_{p}}$, $W_{l}\in\mathbbm{R}^{N_{l}\times N_{l}}$, $W_{up}\in\mathbbm{R}^{N_{u}\times N_{p}}$, $W_{ul}\in\mathbbm{R}^{N_{u}\times N_{l}}$, $W_{pl}\in\mathbbm{R}^{N_{p}\times N_{l}}$. We do not connect any nodes between textual prototypes and few-shot samples because we are only interested in inferring the labels of unlabeled test samples. This results in $W_{p}=\mathbf{0}$, $W_{l}=\mathbf{0}$, $W_{pl}=\mathbf{0}$.
@@
For label propagation, we sparsify $W$ to improve computation cost and robustness. Observations made in Zhu et al. (2023); Kalantidis et al. (2024) suggest that cosine similarities between inputs from different modalities can vary significantly. For example, the cosine similarity between the features of two images is generally significantly lower than between an image feature and a text embedding. Therefore, we search for the K-nearest neighbors within each modality separately as follows, with $\text{NN}k_{\mathcal{D}}$ referring to the $k$ nearest neighbor search within nodes from set $\mathcal{D}$.
@@
Dynamic Graph Expansion: The static graph construction assumes unlabeled test samples are all available at once. This assumption prohibits inductive inference where test samples arrive in a stream fashion. Previous work (Kalantidis et al., 2024) proposed an inductive inference approach by exploiting an external training dataset to build the manifold. With access to such an external dataset, a naive way to support inductive inference is to build a new graph from scratch upon seeing new testing data, which results in high computation complexity. For example, with a total number of $N_{v}=N_{p}+N_{l}+N_{u}$ nodes, each with $d$ dimensions, the computational complexity is $O(d\cdot N_{v}^{3}+N_{v}^{2}\log N_{v})$, which prohibits realistic inference on large scale downstream task. To enable inductive inference while exploiting the testing data manifold, we reduce the complexity to $O(d\cdot N_{v}^{2})$ by a dynamic graph expansion mechanism. Specifically, with an existing adjacency matrix $W$, we use the new test sample $u_{N_{u}+1}$ to query all existing nodes and replace the weakest connections with connections to the new test sample following the update rule in Eq. 5. The update rule will insert a new test sample into the existing graph without re-calculating the K nearest neighbor of the existing nodes.
@@
Incremental Label Propagation: With the dynamically expanded graph, we can perform label propagation incrementally. Specifically, we re-use the pseudo labels from the previous iteration to speed up the convergence of the label propagation process. After inference for each test sample, the pseudo labels are attenuated for label propagation for the next test sample. We follow the rules in Eq. 6 to produce the incremental labels $Y_{u}^{0}$ for the observed test samples for inference on future test samples, where $Y^{T}_{u}$ represents the label predictions obtained after $T$ propagation steps.
@@ 3.4 Context-aware Edge Re-Weighting
The graph construction process depends on a properly chosen distance metric. Measuring the distance between samples based on the features from the vision encoder is subject to the biases learnt by the vision encoder during large-scale pre-training. For example, it is known that VLMs encode features that capture all visual instances in the image, e.g. objects, background, style, etc. For this reason, directly measuring the cosine similarity between the raw features encoded by the VLM is subject to irrelevant information for the downstream task. For example, adapting the VLM to a downstream task of car model classification requires less attention to certain aspects like background or color information. Motivated by this, we propose to re-weight the importance of VLM vision-encoded features for calculating similarity. Specifically, we first calculate the statistics of textual prompts feature dimensions as follows, where $c$ is the feature dimension index.
@@
A higher variance in $c$-th dimension suggests the higher discriminative capability of the dimension, i.e. the more capable it is to discriminate between text prompts. Therefore, we increase the importance of these dimensions with higher variance. Additionally, we calculate the same statistics on the few-shot samples, resulting in $\mu_{c}^{l}$ and $\sigma_{c}^{l}$. In contrast to the context of textual prompts, we argue that high-variance feature dimensions in the context of few-shot samples highlight the intra-class variance and that this should be suppressed when calibrating the similarity. Therefore, we use the reciprocal of the variance for re-weighting graph edge similarity. The new edge weights are presented in Eq. 8: